OpenStudy (anonymous):

An X-ray tube is operated at 10kV.

OpenStudy (kropot72):

The velocity v of an electron striking the anticathode can be found from: \[v=\sqrt{\frac{2eV}{m}}\] where e is the charge on the electron = 1.6 * 10^-19 C V is the accelerating voltage = 10^4 V m is the mass of an electron = 9 * 10^-31 kg Plugging in these values we get: \[v=\sqrt{\frac{2\times1.6\times10^{-19} \times10^{4}}{9\times10^{-31}}}=you\ can\ calculate\]